OpenStudy (anonymous):

∫dx/x^2√1-x^2 let,x=sinz =>dx/dz=d/dz sinz =>dx/dz=cosz =>dx=cosz dz Now,∫dx/x^2√1-x^2=∫cosz dz/sin^2z√1-sin^2z =∫cosz dz/sin^2z√cos^2z=∫cosz dz/sin^2z.cosz=∫dz/sin^2z=∫dz/1/cosec^2z=∫dz cosec^2z=-cotz+c then what? @ganeshie8

ganeshie8 (ganeshie8):

\(\large \mathbb {\int \frac{dx}{x^2\sqrt{1-x^2}} }\)

ganeshie8 (ganeshie8):

its like that ?

OpenStudy (anonymous):

yes ! :)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

please help..i really need help :(

ganeshie8 (ganeshie8):

so far looks good, u need to given answer in x's.. not z..

OpenStudy (anonymous):

yeah its like -cotz=-cosz/sinz=-cosz/x

ganeshie8 (ganeshie8):

yes, wat u gona do wid numerator cosz ?

OpenStudy (anonymous):

there cosz=dz/dx

OpenStudy (anonymous):

if you see up

OpenStudy (anonymous):

sorry dx/dz=cosz

OpenStudy (anonymous):

dx/dz/x=dx/dz.1/x

ganeshie8 (ganeshie8):

-cotz+c your answer is right so far. u just need to get rid of z's ok

ganeshie8 (ganeshie8):

let me show u quick how to get rid of z :)

OpenStudy (anonymous):

:) okay

OpenStudy (nincompoop):

I find this interesting

ganeshie8 (ganeshie8):

u have x = sin z draw the right triangle representing this..

ganeshie8 (ganeshie8):